A store selects 4 items from a selection of 6 items to arrange in a display. How many different arrangements are possible?

Is my solution of 6 * 4 = 24 a correct answer or to easy???

Answer by stanbon(48558)

A store selects 4 items from a selection of 6 items to arrange in a display. How many different arrangements are possible?
---------
Comment: Arrangements are called permutations.
You are selecting 4 of 6.
Your Problem:
6P4 = [6/!]/[(6-4)!] = 6*5*4*3 = 360 arrangements
========
Cheers,
Stan H.